OpenStudy (anonymous):

A coat costs $159.99 and is on clearance for 50% off. When you get to registered you get an additional 20% off the discounted price. To the nearest cent, how much will the coat cost?

OpenStudy (anonymous):

50% off means half the cost off so left is 79.995 additional 20% off = 20/100* 79.995 = 0.20*79.995 = 15.999 so final cost = 79.995 - 15.999 = 63.996 = 64 $